how do you solve these 2 questions: 3/4c=12 and 11/12=3/4h

Answer by checkley71(8403)

3c/4=12
cross multiply
3c=4*12
3c=48
c=48/3
c=16 answer.
-----------------------------------------
IF THE EQUATION IS: 11/12=3/(4H)
THEN:
36=44H
H=36/44
H=9/11
------------------------------------------
I INTERPRETED THIS EQUATION THE SAME AS THE FIRST EQUATION. THAT THE UNKNOWN WAS A MULTIPLIER OF THE FRACTON - NOT PART OF THE DENOMINATOR.
11/12=3h/4
3h*12=11*4
36h=44
h=44/36
h=11/9 answer.